the avg velocity of a body moving with uniform acc. after travelling a distance of 3.06 m is 0.34 m/s. If the change in velocity of the body is 0.8 m/s during this time its uniform acc. is A : 0.01B : 0.02C : 0.03D : 0.04

Grade:12

2 Answers

Sandeep Pathak
9 years ago
Let the initial velocity by u and final velocity be v. Then,

Average velocity,<v> = \frac{v+u}{2} = 0.34 m/s
Change in velocity isv-u = 0.8m/s
This gives,v = 0.74m/s, u = -0.06m/s
Usingv^2=u^2+2as, you can find the answer.
Yokesh K.K.
5 years ago
Average velocity = Total displacement/ Total time
Since it is a straight line motion , displacement = distance. 
Therefore, we get t = 9 s
Now
V-U = 2aS
Therefore , a = 0.02 m/s ^2
Correct me if I'm wrong